# A Survey of Large Language Models for Graphs

### v1

@@ 1. Introduction
Graphs, comprising nodes and edges that signify relationships, are essential for illustrating real-world connections across various domains. These include social networks (Myers et al., 2014; Li et al., 2024b), molecular graphs (Cao et al., 2023), recommender systems (He et al., 2020), and academic networks (Hu et al., 2020). This structured data form is integral in mapping complex interconnections relevant to a wide range of applications.
@@
In recent years, Graph Neural Networks (GNNs) have excelled in tasks like node classification and link prediction by passing and aggregating information across nodes and iteratively refining node features through supervised learning. They use graph labels to enhance model accuracy and capture structural nuances. Notable models include Graph Convolutional Networks (GCNs)(Kipf and Welling, 2016) for effective embedding propagation, Graph Attention Networks (GAT)(Velickovic et al., 2017) that leverage attention mechanisms for precise aggregation, and Graph Transformers (Yun et al., 2019) which employ self-attention and positional encoding for global insights. Innovations such as Nodeformer (Wu et al., 2022) and DIFFormer (Wu et al., 2023) address scalability in large graphs with efficient attention and differentiable pooling techniques. Despite these advancements, current GNN methodologies still confront challenges such as data sparsity (You et al., 2021) and limited generalizability (Garg et al., 2020; Xia et al., 2024).
@@
Large Language Models (LLMs) have demonstrated exceptional language understanding and summarization capabilities in Natural Language Processing, showcasing strong abilities for previously unseen tasks (Devlin et al., 2018; Raffel et al., 2020). Their advent has sparked significant interest within the graph learning community, prompting investigations into the potential of LLMs to enhance performance on graph-related tasks. Approaches such as InstructGLM (Ye et al., 2023) and NLGraph (Wang et al., 2024a) have developed prompts that enable LLMs to understand graph structures and respond to queries effectively. Alternatively, methods like GraphGPT (Tang et al., 2023) and GraphLLM (Chai et al., 2023) integrate GNNs to feed tokens into the LLMs, allowing them to understand graph structures. This synergy has not only improved task performance but also demonstrated impressive zero-shot generalization capabilities.
@@
We offer a comprehensive review of the current state-of-the-art Large Language Models (LLMs) for graph learning, elucidating their strengths and pinpointing their limitations.
We introduce a novel taxonomy for categorizing existing research based on their framework design, which provides a deeper insight into how LLMs can be seamlessly integrated with graph learning.
We explore potential avenues for future research, including addressing the prevalent challenges in merging LLMs with graph learning methods and venturing into novel application areas.
@@ 2.1. Definitions
Graph-Structured Data. In computer science, a graph is a non-linear data structure that consists of a set of nodes , and a set of edges connecting these nodes. Each edge is associated with a pair of nodes , where and are the endpoints of the edge. The edge may be directed, meaning it has a orientation from to , or undirected, meaning it has no orientation. Furthermore, A Text-attributed Graph (TAG) is a graph that assigns a sequential text feature (sentence) to each node, denoted as , which is widely used in the era of large language models. The TAG can be formally represented as , where is the set of text features.
@@
Large Language Models (LLMs). Language Models (LMs) is a statistical model that estimate the probability distribution of words for a given sentence. Recent research has shown that LMs with billions of parameters exhibit superior performance in solving a wide range of natural language tasks (e.g., translation, summarization and instruction following), making them large language models (LLMs). In general, most recent LLMs are built with transformer blocks that use a query-key-value (QKV)-based attention mechanism to aggregate information in the sequence of tokens. Based on the direction of attention, large language models (LLMs) can be categorized into two types (given a sequence of tokens ):
Causal Language Modeling (CLM). The model only considers the left context to predict the next token:
@@ 2.2. Taxonomy
LLMs-Graphs Intergration. In this line, LLMs achieve a higher level of integration with graph data, such as fusion training or alignment with GNNs, and also build LLM-based agents to interact with graph information.
@@ 3.1.1. Node-level Tokenization
Within this line, GraphGPT (Tang et al., 2023) proposes to initially align the graph encoder with natural language semantics through text-graph grounding, and then combine the trained graph encoder with the LLM using a projector. Through the two-stage instruction tuning paradigm, the model can directly complete various graph learning downstream tasks with natural language, thus perform strong zero-shot transferability and multi-task compatibility. The proposed Chain-of-Thought distillation method empowers GraphGPT to migrate to complex tasks with small parameter sizes. Then, HiGPT (Tang et al., 2024) proposes to combine the language-enhanced in-context heterogeneous graph tokenizer with LLMs, solving the challenge of relation type heterogeneity shift between different heterogeneous graphs. Meanwhile, the two-stage heterogeneous graph instruction-tuning injects both homogeneity and heterogeneity awareness into the LLM. And the Mixture-of-Thought (MoT) method combined with various prompt engineering further solves the common data scarcity problem in heterogeneous graph learning. GIMLET (Zhao et al., 2024), as a unified graph-text model, leverages natural language instructions to address the label insufficiency challenge in molecule-related tasks, effectively alleviating the reliance on expensive lab experiments for data annotation. It employs a generalized position embedding and attention mechanism to encode both graph structures and textual instructions as a unified token combination that is fed into a transformer decoder. GraphTranslator (Zhang et al., 2024) proposes the use of a translator with shared self-attention to align both the target node and instruction, and employs cross attention to map the node representation encoded by the graph model to fixed-length semantic tokens. The proposed daul-phase training paradigm empowers the LLM to make predictions based on language instructions, providing a unified solution for both pre-defined and open-ended graph-based tasks. Instead of using pre-computed node features of varying dimensions, UniGraph (He and Hooi, 2024) leverages Text-Attributed Graphs for unifying node representations, featuring a cascaded architecture of language models and graph neural networks as backbone networks.
@@ 3.1.2. Graph-level Tokenization
Within this domain, GraphLLM (Chai et al., 2023) utilizes a graph transformer that incorporates the learnable query and positional encoding to encode the graph structure and obtain graph representations through pooling. These representations are directly used as graph-enhanced prefix for prefix tuning in the LLM, demonstrating remarkable effectiveness in fundamental graph reasoning tasks. MolCA (Liu et al., 2023b) with Cross-Modal Projector and Uni-Modal Adapter is a method that enables a language model to understand both text- and graph-based molecular contents through the proposed dual-stage pre-training and fine-tuning stage. It employs a cross-modal projector implemented as a Q-Former to connect a graph encoder’s representation space and a language model’s text space, and a uni-modal adapter for efficient adaptation to downstream tasks. InstructMol (Cao et al., 2023) introduces a projector that aligns the molecular graph encoded by the graph encoder with the molecule’s Sequential information and natural language instructions, with the first stage of Alignment Pretraining and the second stage of Task-specific Instruction Tuning enabling the model to achieve excellent performance in various drug discovery-related molecular tasks. GIT-Mol (Liu et al., 2024b) further unifies the graph, text, and image modalities through interaction cross-attention between different modality encoders, and aligns these three modalities, enabling the model to simultaneously perform four downstream tasks: captioning, generation, recognition, and prediction. GNP (Tian et al., 2024) employs cross-modality pooling to integrate the node representations encoded by the graph encoder with the natural language tokens, resulting in a unified graph representation. This representation is aligned with the instruction through the LLM to demonstrate superiority in commonsense and biomedical reasoning tasks. Recently, G-Retriever (He et al., 2024b) utilizes retrieval-augmented techniques to obtain subgraph structures. It completes various downstream tasks in GraphQA (Graph Question Answering) through the collaboration of graph encoder and LLMs.
@@ 3.2.1. Embeddings from LLMs for GNNs
In this domain, G-Prompt (Huang et al., 2023) adds a GNN layer at the end of a pre-trained language models (PLMs) to achieve graph-aware fill-masking self-supervised learning. By doing so, G-Prompt can generate task-specific, explainable node embeddings for downstream tasks using prompt tuning. SimTeG (Duan et al., 2023) first leverages parameter-efficient fine-tuning on the text embeddings obtained by LLMs for downstream tasks (e.g., node classification). Then, the node embeddings are fed into GNNs for inference. Similarly, GALM (Xie et al., 2023) utilizes BERT as a pre-language model to encode text embeddings for each node. Then, the model is pre-trained through unsupervised learning tasks, such as link prediction, to minimize empirical loss and find optimal model parameters, which enables GALM to be applied for various downstream tasks. Recently, OFA (Liu et al., 2024a) leverages LLMs to unify graph data from different domains into a common embedding space for cross-domain learning. It also uses LLMs to encode task-relevant text descriptions for constructing prompt graphs, allowing the model to perform specific tasks based on context. TAPE (He et al., 2024a) uses customized prompts to query LLMs, generating both prediction and text explanation for each node. Then, DeBERTa is fine-tuned to convert the text explanations into node embeddings for GNNs. Finally, GNNs can use a combination of the original text features, explanation features, and prediction features to predict node labels. In the field of recommendation, LLMRec (Wei et al., 2024) achieves graph augmentation on user-item interaction data using GPT-3.5, which not only filters out noise interactions and adds meaningful training data, but also enriches the initial node embeddings for users and items with generated rich textual profiles, ultimately improving the performance of recommenders.
@@ 3.2.2. Labels from LLMs for GNNs
Follow this line, OpenGraph (Xia et al., 2024) employs LLMs to generate nodes and edges, mitigating the issue of sparse training data. The generation process for nodes and edges is refined using the Gibbs sampling algorithm and a tree-of-prompt strategy, which is then utilized to train the graph foundation model. LLM-GNN (Chen et al., 2023) leverages LLMs as annotators to generate node category predictions with confidence scores, which serve as labels. Post-filtering is then employed to filter out low-quality annotations while maintaining label diversity. Finally, the generated labels are used to train GNNs. GraphEdit (Guo et al., 2024) leverages the LLMs to build an edge predictor, which is used to evaluate and refine candidate edges against the original graph’s edges. In recommender systems, RLMRec (Ren et al., 2023) leverages LLMs to generate text descriptions of user/item preferences. These descriptions are then encoded as semantic embeddings to guide the representation learning of ID-based recommenders using contrastive and generative learning techniques.
@@ 3.2.3. Discussion
3.3. LLMs-Graphs Intergration
@@ 3.3.1. Alignment between GNNs and LLMs
Within this topic, MoMu (Su et al., 2022) is a multimodal molecular foundation model that includes two separate encoders, one for handling molecular graphs (GIN) and another for handling text data (BERT). It uses contrastive learning to pre-train the model on a dataset of molecular graph-text pairs. This approach enables MoMu to directly imagine new molecules from textual descriptions. Also in the bioinfo domain, MoleculeSTM (Liu et al., 2023a) combines the chemical structure information of molecules (i.e., molecular graph) with their textual descriptions (i.e., SMILES strings), and uses a contrastive learning to jointly learn the molecular structure and textual descriptions. It show great performance on multiple benchmark tests, including structure-text retrieval, text-based editing tasks, and molecular property prediction. Similarly, in ConGraT (Brannon et al., 2023), a contrastive graph-text pretraining technique is proposed to align the node embeddings encoded by LMs and GNNs simultaneously. The experiments are conducted on social networks, citation networks, and link networks, and show great performance on node and text classification as well as link prediction tasks. Furthermore, G2P2 (Wen and Fang, 2023) enhances graph-grounded contrastive pre-training by proposing three different types of alignment: text-node, text-summary, and node-summary alignment. This enables G2P2 to leverage the rich semantic relationships in the graph structure to improve text classification performance in low-resource environments. GRENADE (Li et al., 2023a) is a graph-centric language model that proposes graph-centric contrastive learning and knowledge alignment to achieve both node-level and neighborhood-level alignment based on the node embeddings encoded from GNNs and LMs. This enables the model to capture text semantics and graph structure information through self-supervised learning, even in the absence of human-annotated labels. In addition to contrastive learning, THLM (Zou et al., 2023) leverages BERT and HGNNs to encode node embeddings and uses a positive-negative classification task with negative sampling to improve the alignment of embeddings from two different modalities. Recently, GLEM (Zhao et al., 2022) adopts an efficient and effective solution that integrates graph structure and language learning through a variational expectation-maximization (EM) framework. By iteratively using LMs and GNNs to provide labels for each other in node classification, GLEM aligns their capabilities in graph tasks.

### v3

@@ 1. Introduction
Graphs, comprising nodes and edges that signify relationships, are essential for illustrating real-world connections across various domains. These include social networks (Myers et al., 2014; Li et al., 2024b), molecular graphs (Cao et al., 2023), recommender systems (He et al., 2020; Ren et al., 2023), and academic networks (Hu et al., 2020). This structured data form is integral in mapping complex interconnections relevant to a wide range of applications.
@@
In recent years, Graph Neural Networks (GNNs) (Wu et al., 2020) have emerged as a powerful tool for a variety of tasks, including node classification (Xu et al., 2018) and link prediction (Zhang and Chen, 2018). By passing and aggregating information across nodes and iteratively refining node features through supervised learning, GNNs have achieved remarkable results in capturing structural nuances and enhancing model accuracy. To accomplish this, GNNs leverage graph labels to guide the learning process. Several notable models have been proposed in the literature, each with its own strengths and contributions. For instance, Graph Convolutional Networks (GCNs) (Kipf and Welling, 2016) have been shown to be effective in propagating embeddings across nodes, while Graph Attention Networks (GATs) (Veličković et al., 2017) leverage attention mechanisms to perform precise aggregation of node features. Additionally, Graph Transformers (Yun et al., 2019; Li et al., 2023c) employ self-attention and positional encoding to capture global signals among the graph, further improving the expressiveness of GNNs. To address scalability challenges in large graphs, methods such as Nodeformer (Wu et al., 2022) and DIFFormer (Wu et al., 2023b) have been proposed. These approaches employ efficient attention mechanisms and differentiable pooling techniques to reduce computational complexity while maintaining high levels of accuracy. Despite these advancements, current GNN methodologies still face several challenges. For example, data sparsity remains a significant issue, particularly in scenarios where the graph structure is incomplete or noisy (You et al., 2021). Moreover, the generalization ability of GNNs to new graphs or unseen nodes remains an open research question, with recent works highlighting the need for more robust and adaptive models (Garg et al., 2020; Xia et al., 2024; Zhao et al., 2024b).
@@
Large Language Models (LLMs) (Zhao et al., 2023a), which show great generalization abilities for unseen tasks (Devlin et al., 2018; Raffel et al., 2020; Wang et al., 2023b), have emerged as powerful tools in various research fields, including natural language processing (Achiam et al., 2023), computer vision (Liu et al., 2024b, a), and information retrieval (Zhu et al., 2023; Hou et al., 2024; Lin et al., 2024). The advent of LLMs has sparked significant interest within the graph learning community (Huang et al., 2024b; Jin et al., 2023; Li et al., 2023b), prompting investigations into the potential of LLMs to enhance performance on graph-related tasks. Researchers have explored various approaches to leverage the strengths of LLMs for graph learning, resulting in a new wave of methods that combine the power of LLMs with graph neural networks. One promising direction is to develop prompts that enable LLMs to understand graph structures and respond to queries effectively. For instance, approaches such as InstructGLM (Ye et al., 2023) and NLGraph (Wang et al., 2024a) have designed specialized prompts that allow LLMs to reason over graph data and generate accurate responses. Alternatively, other methods have integrated GNNs to feed tokens into the LLMs, allowing them to understand graph structures more directly. For example, GraphGPT (Tang et al., 2023) and GraphLLM (Chai et al., 2023) use GNNs to encode graph data into tokens, which are then fed into the LLMs for further processing. This synergy between LLMs and GNNs has not only improved task performance but also demonstrated impressive zero-shot generalization capabilities, where the models can accurately answer queries about unseen graphs or nodes.
@@
Comprehensive Review of LLMs for Graph Learning. We offer a comprehensive review of the current state-of-the-art Large Language Models (LLMs) for graph learning, elucidating their strengths and pinpointing their limitations.
Novel Taxonomy for Categorizing Research. We introduce a novel taxonomy for categorizing existing research based on their framework design, which provides a deeper insight into how LLMs can be seamlessly integrated with graph learning.
Future Research Avenues. We also explore potential avenues for future research, including addressing the prevalent challenges in merging LLMs with graph learning methods and venturing into novel application areas.
@@ 2.1. Definitions
Graph-Structured Data. In computer science, a graph is a non-linear data structure that consists of a set of nodes , and a set of edges connecting these nodes. Each edge is associated with a pair of nodes , where and are the endpoints of the edge. The edge may be directed, meaning it has a orientation from to , or undirected, meaning it has no orientation. Furthermore, A Text-Attributed Graph (TAG) is a graph that assigns a sequential text feature (i.e., sentence) to each node, denoted as , which is widely used in the era of large language models. The text-attributed graph can be formally represented as , where is the set of text features.
@@
Large Language Models (LLMs). Language Models (LMs) is a statistical model that estimate the probability distribution of words for a given sentence. Recent research has shown that LMs with billions of parameters exhibit superior performance in solving a wide range of natural language tasks (e.g., translation, summarization and instruction following), making them Large Language Models (LLMs). In general, most recent LLMs are built with transformer blocks that use a query-key-value (QKV)-based attention mechanism to aggregate information in the sequence of tokens. Based on the direction of attention, LLMs can be categorized into two types (given a sequence of tokens ):
Masked Language Modeling (MLM). Masked Language Modeling is a popular pre-training objective for LLMs that involves masking out certain tokens in a sequence and training the model to predict the masked tokens based on the surrounding context. Specifically, the model takes into account both the left and right context of the masked token to make accurate predictions:
(2) Representative models include BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019).
Causal Language Modeling (CLM). Causal Language Modeling is another popular training objective for LLMs that involves predicting the next token in a sequence based on the previous tokens. Specifically, the model only considers the left context of the current token to make accurate predictions:
@@ 2.2. Taxonomy
LLMs-Graphs Integration. In this line, LLMs achieve a higher level of integration with graph data, such as fusion training or alignment with GNNs, and also build LLM-based agents to interact with graph information.
@@ 3.1.1. Node-level Tokenization
Within this line, GraphGPT (Tang et al., 2023) proposes to initially align the graph encoder with natural language semantics through text-graph grounding, and then combine the trained graph encoder with the LLM using a projector. Through the two-stage instruction tuning paradigm, the model can directly complete various graph learning downstream tasks with natural language, thus perform strong zero-shot transferability and multi-task compatibility. The proposed Chain-of-Thought distillation method empowers GraphGPT to migrate to complex tasks with small parameter sizes. Then, HiGPT (Tang et al., 2024) proposes to combine the language-enhanced in-context heterogeneous graph tokenizer with LLMs, solving the challenge of relation type heterogeneity shift between different heterogeneous graphs. Meanwhile, the two-stage heterogeneous graph instruction-tuning injects both homogeneity and heterogeneity awareness into the LLM. And the Mixture-of-Thought (MoT) method combined with various prompt engineering further solves the common data scarcity problem in heterogeneous graph learning. GIMLET (Zhao et al., 2024a), as a unified graph-text model, leverages natural language instructions to address the label insufficiency challenge in molecule-related tasks, effectively alleviating the reliance on expensive lab experiments for data annotation. It employs a generalized position embedding and attention mechanism to encode both graph structures and textual instructions as a unified token combination that is fed into a transformer decoder. GraphTranslator (Zhang et al., 2024) proposes the use of a translator with shared self-attention to align both the target node and instruction, and employs cross attention to map the node representation encoded by the graph model to fixed-length semantic tokens. The proposed daul-phase training paradigm empowers the LLM to make predictions based on language instructions, providing a unified solution for both pre-defined and open-ended graph-based tasks. Instead of using pre-computed node features of varying dimensions, UniGraph (He and Hooi, 2024) leverages Text-Attributed Graphs for unifying node representations, featuring a cascaded architecture of language models and graph neural networks as backbone networks. In recent research on recommendation systems, XRec (Ma et al., 2024) has been proposed as a method that utilizes the encoded user/item embeddings from graph neural networks as collaborative signals. These signals are then integrated into each layer of large language models, enabling the generation of explanations for recommendations, even in zero-shot scenarios.
@@ 3.1.2. Graph-level Tokenization
Within this domain, GraphLLM (Chai et al., 2023) utilizes a graph transformer that incorporates the learnable query and positional encoding to encode the graph structure and obtain graph representations through pooling. These representations are directly used as graph-enhanced prefix for prefix tuning in the LLM, demonstrating remarkable effectiveness in fundamental graph reasoning tasks. MolCA (Liu et al., 2023b) with Cross-Modal Projector and Uni-Modal Adapter is a method that enables a language model to understand both text- and graph-based molecular contents through the proposed dual-stage pre-training and fine-tuning stage. It employs a cross-modal projector implemented as a Q-Former to connect a graph encoder’s representation space and a language model’s text space, and a uni-modal adapter for efficient adaptation to downstream tasks. InstructMol (Cao et al., 2023) introduces a projector that aligns the molecular graph encoded by the graph encoder with the molecule’s Sequential information and natural language instructions, with the first stage of Alignment Pretraining and the second stage of Task-specific Instruction Tuning enabling the model to achieve excellent performance in various drug discovery-related molecular tasks. GIT-Mol (Liu et al., 2024c) further unifies the graph, text, and image modalities through interaction cross-attention between different modality encoders, and aligns these three modalities, enabling the model to simultaneously perform four downstream tasks: captioning, generation, recognition, and prediction. GNP (Tian et al., 2024) employs cross-modality pooling to integrate the node representations encoded by the graph encoder with the natural language tokens, resulting in a unified graph representation. This representation is aligned with the instruction through the LLM to demonstrate superiority in commonsense and biomedical reasoning tasks. Recently, G-Retriever (He et al., 2024) utilizes retrieval-augmented techniques to obtain subgraph structures. It completes various downstream tasks in GraphQA (Graph Question Answering) through the collaboration of graph encoder and LLMs.
@@ 3.2.1. Embeddings from LLMs for GNNs
In this domain, G-Prompt (Huang et al., 2023) adds a GNN layer at the end of a pre-trained language models (PLMs) to achieve graph-aware fill-masking self-supervised learning. By doing so, G-Prompt can generate task-specific, explainable node embeddings for downstream tasks using prompt tuning. SimTeG (Duan et al., 2023) first leverages parameter-efficient fine-tuning on the text embeddings obtained by LLMs for downstream tasks (e.g., node classification). Then, the node embeddings are fed into GNNs for inference. Similarly, GALM (Xie et al., 2023) utilizes BERT as a pre-language model to encode text embeddings for each node. Then, the model is pre-trained through unsupervised learning tasks, such as link prediction, to minimize empirical loss and find optimal model parameters, which enables GALM to be applied for various downstream tasks. Recently, OFA (Liu et al., 2023c) leverages LLMs to unify graph data from different domains into a common embedding space for cross-domain learning. It also uses LLMs to encode task-relevant text descriptions for constructing prompt graphs, allowing the model to perform specific tasks based on context. TAPE (He et al., 2023) uses customized prompts to query LLMs, generating both prediction and text explanation for each node. Then, DeBERTa is fine-tuned to convert the text explanations into node embeddings for GNNs. Finally, GNNs can use a combination of the original text features, explanation features, and prediction features to predict node labels. In the field of recommendation, LLMRec (Wei et al., 2024) achieves graph augmentation on user-item interaction data using GPT-3.5, which not only filters out noise interactions and adds meaningful training data, but also enriches the initial node embeddings for users and items with generated rich textual profiles, ultimately improving the performance of recommenders.
@@ 3.2.2. Labels from LLMs for GNNs
Follow this line, OpenGraph (Xia et al., 2024) employs LLMs to generate nodes and edges, mitigating the issue of sparse training data. The generation process for nodes and edges is refined using the Gibbs sampling algorithm and a tree-of-prompt strategy, which is then utilized to train the graph foundation model. LLM-GNN (Chen et al., 2023) leverages LLMs as annotators to generate node category predictions with confidence scores, which serve as labels. Post-filtering is then employed to filter out low-quality annotations while maintaining label diversity. Finally, the generated labels are used to train GNNs. GraphEdit (Guo et al., 2024) leverages the LLMs to build an edge predictor, which is used to evaluate and refine candidate edges against the original graph’s edges. In recommender systems, RLMRec (Ren et al., 2024b) leverages LLMs to generate text descriptions of user/item preferences. These descriptions are then encoded as semantic embeddings to guide the representation learning of ID-based recommenders using contrastive and generative learning techniques (Ren et al., 2024a).
@@ 3.2.3. Discussion
3.3. LLMs-Graphs Integration
@@ 3.3.1. Alignment between GNNs and LLMs
Within this topic, MoMu (Su et al., 2022) is a multimodal molecular foundation model that includes two separate encoders, one for handling molecular graphs (GIN) and another for handling text data (BERT). It uses contrastive learning to pre-train the model on a dataset of molecular graph-text pairs. This approach enables MoMu to directly imagine new molecules from textual descriptions. Also in the bioinfo domain, MoleculeSTM (Liu et al., 2023a) combines the chemical structure information of molecules (i.e., molecular graph) with their textual descriptions (i.e., SMILES strings), and uses a contrastive learning to jointly learn the molecular structure and textual descriptions. It show great performance on multiple benchmark tests, including structure-text retrieval, text-based editing tasks, and molecular property prediction. Similarly, in ConGraT (Brannon et al., 2023), a contrastive graph-text pretraining technique is proposed to align the node embeddings encoded by LMs and GNNs simultaneously. The experiments are conducted on social networks, citation networks, and link networks, and show great performance on node and text classification as well as link prediction tasks. Furthermore, G2P2 (Wen and Fang, 2023a, b) enhances graph-grounded contrastive pre-training by proposing three different types of alignment: text-node, text-summary, and node-summary alignment. This enables G2P2 to leverage the rich semantic relationships in the graph structure to improve text classification performance in low-resource environments. GRENADE (Li et al., 2023a) is a graph-centric language model that proposes graph-centric contrastive learning and knowledge alignment to achieve both node-level and neighborhood-level alignment based on the node embeddings encoded from GNNs and LMs. This enables the model to capture text semantics and graph structure information through self-supervised learning, even in the absence of human-annotated labels. In addition to contrastive learning, THLM (Zou et al., 2023) leverages BERT and HGNNs to encode node embeddings and uses a positive-negative classification task with negative sampling to improve the alignment of embeddings from two different modalities. Recently, GLEM (Zhao et al., 2022) adopts an efficient and effective solution that integrates graph structure and language learning through a variational expectation-maximization (EM) framework. By iteratively using LMs and GNNs to provide labels for each other in node classification, GLEM aligns their capabilities in graph tasks.